1. A physician ordered Magnesium Sulfate 40 grams in 1,000mL of LR at a rate of 2g/hour for a woman in preterm labor. What rate should the infusion pump be set at?

2. Betony is a 21/2 year old who is newly diagnosed with HIV infection. She is to be started on a Retrovir 120 mg by mouth every 12 hours. She weighs 13 kg and is 90 cm long.

  1. Using the formula for BSA calculate her body surface area: BSA (m2) = Ht (cm) x wt (kg) / 3600
  2. Retrovir is prescribed on an individual basis ranging from 180 to 240 mg/m2 per dose every 12 hours. It comes in a concentration of 10 mg/mL.

Is her does of 120 mg within the acceptable range?

How many mL will you administer?

3. The order is for Pitocin 2 mu/min for a woman undergoing induction of labor. Pitocin is available in vials of 10 units/mL. The protocol on the Labor and Delivery unit is to mix 2 vials in 1,000mL of NS. What should the nurse set the IV pump at?
Answer #1

1 Mgso4 hogm/100oml 24 Housn Desired dose Thour. 29 40gm/1000ml 19m/100g ml - 19 m /a5 me 2gm /houn 50 ml /Hour Dei red doseBetony- P1 /2 year BSAGM) = Ht com) x ut(kg)/3600 =go x 13 3600 70/3600 1170 = 3.2 (m³) BSA 18070240 mgm. * Retaoir IS preica@3 Desired desse = &mu /min. 1 val contaiùng 10Units /ml 1000ml 2 viald în. Mix 1000 ml 10X 2=2OU nits 9ou 1 ml containing =